OpenStudy (anonymous):

An experiment consists of rolling a fair die and flipping a fair coin. What is the probability of getting an even number and a tail? 1/2 correct?

OpenStudy (shubhamsrg):

3/6 * 1/2

OpenStudy (shubhamsrg):

=1/4

OpenStudy (chaise):

3/6 * 1/2 = 1/6 There is a three in six chance of getting a even number on a dice with numbers 1-6, and a 1 in 2 chance of getting a tail on a coin. Multiply the two chances together.

OpenStudy (anonymous):

3/6 from where??

OpenStudy (anonymous):

nvm thx